OpenStudy (anonymous):

six people are going to sit in a row on a bench. romeo wants to sit next to juliet. caesar wants to sit next to brutus. homer and pierre can sit anywhere. how many ways can these six people be seated?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

pack romeo and juliet in one bag pack caeser and brutus in another bag leave homer and pierre free

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

So you have `two bags` and `two people` to arrange

ganeshie8 (ganeshie8):

that means `four` objects to arrange

ganeshie8 (ganeshie8):

How many ways can you arrange `four` objects ?

OpenStudy (anonymous):

8??

ganeshie8 (ganeshie8):

four objects can be arranged in 4! ways : 4 ways for choosing first position, after that, 3 ways for choosing second position, after that, 2 ways for choosing third position, after that, 1 way for choosing last position so total number of ways for arranging 4 objects = 4*3*2*1 = 4! = 24

ganeshie8 (ganeshie8):

@dinisha

ganeshie8 (ganeshie8):

we're not done yet - as you still can arrange the people inside bags : 2! ways for arranging people in first bag 2! ways for arranging people in second bag

ganeshie8 (ganeshie8):

So, total number of ways of arranging these 6 people honoring given requirements is : 4! * 2! * 2! = 24*2*2 = 96

ganeshie8 (ganeshie8):

if that makes any sense...

OpenStudy (anonymous):

ya!! that you..